# OG-VLA: 3D-Aware Vision Language Action Model via Orthographic Image Generation

### v1

@@ 4 Simulation Experiments
OG-VLA@30k 86.7$\pm$2.9 15.0$\pm$8.7 38.3$\pm$2.9 51.7$\pm$2.9 0.0$\pm$0.0 16.7$\pm$2.9 25.0$\pm$5.0 16.7$\pm$7.6 31.2$\pm$2.9
   -Novel Object 85.0$\pm$5.0 0.0$\pm$0.0 1.7$\pm$2.9 55.0$\pm$13.2 1.7$\pm$2.9 5.0$\pm$5.0 18.3$\pm$2.9 6.7$\pm$7.6 21.7$\pm$0.7
   -Novel Scene 70.0$\pm$2.9 1.7$\pm$2.8 26.7$\pm$11.5 36.7$\pm$5.8 1.7$\pm$2.9 1.7$\pm$2.9 16.7$\pm$11.5 8.3$\pm$2.9 20.8$\pm$1.3
   -Novel State 0.0$\pm$0.0 13.3$\pm$7.6 13.3$\pm$2.9 20.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 8.3$\pm$7.6 13.3$\pm$2.9 8.5$\pm$1.9
@@
Table 1: Success rate on ARNOLD [12]. Success rates for 8 tasks and 4 test splits are shown for 2 baseline models and our model at specified training iterations (30k and 100k). The first row for each model is the Novel Pose split (the Test set in ARNOLD). The numbers in bold represent the best system for each task and test split.
4.1 Benchmarks and Datasets
We evaluate our method on the Arnold [12] and Colosseum [11] benchmarks, which test language-grounded robot task learning in realistic 3D scenes with emphasis on testing generalization. Arnold uses five input cameras (front, base, left, wrist top, wrist bottom) and includes eight language-conditioned tasks (see Table 1), each with four generalization test splits: (1) Novel Pose (held-out object/robot placements), (2) Novel Object (unseen objects), (3) Novel Scene (unseen scenes with seen objects), and (4) Novel State (unseen goal states). Arnold tasks follow a two-keyframe format—grasping and manipulating (e.g., pull drawer to 50% open)—and do not require gripper state prediction. These tasks demand both object pose and free space reasoning. Colosseum has four cameras (front, left shoulder, right shoulder, wrist) and features 20 language-conditioned tabletop tasks (e.g., close box, empty dishwasher) with 2–13 gripper keyframes (average 6), requiring gripper open/close prediction. It evaluates generalization via the all perturbation test set, which simultaneously alters object/table/background appearance, lighting, camera pose, and adds distractors.
@@ 4.1 Benchmarks and Datasets
Each Arnold test split includes 20 episodes; Colosseum uses 25. We evaluate each Arnold model over 3 runs to account for simulator’s motion planner noise during keypoint execution, and we report means and standard deviations. Baseline models were trained by the respective benchmark authors. We re-run PerAct [7] for 3 runs on Arnold for fair comparison, and report 6D-CLIPort [38] results from Arnold [12] due to lack of checkpoint release.
4.2 Simulation Results
Arnold
@@ 4.2 Simulation Results
We present our model results in Table 1. OG-VLA outperforms the baseline (PerAct) for most of the tasks’ Novel Pose split (seen objects and scenes), with a task-averaged relative improvement of 10.8%. We present our result at 30k and 100k iterations. OG-VLA improves significantly across generalization splits (Novel Object, Scene, and State), with relative increases of 20.0% and 46.5% respectively at 30k and 100k iterations in overall success rates. This demonstrates the effectiveness of our method in adapting pretrained visual and textual priors to robotics applications.
@@
Figure 3: Evaluation on the COLOSSEUM benchmark [11]. Task-averaged success rate shows that OG-VLA outperforms all baselines on the hardest generalization test set (all perturbation).
Colosseum
@@
Figure 3 shows task-averaged success rate on all perturbation test set on the Colosseum benchmark. We compare with baselines reported in Colosseum, including R3M [19], MVP [20], 3DDA [39], RVT [8], and PerAct [7]. OG-VLA improves upon the baselines by a relative increase of 45.8%. The absolute performance remains low however (10.5%). We observe that Colosseum tasks are much harder to learn due to longer sequence of keyframe predictions, leading to error accumulation for an imitation learning based method.
4.3 Design Choice Ablations
@@
We conducted extensive ablations on the Arnold benchmark to validate our design choices. We highlight the main findings in this section, with detailed results and analysis in Appendix C.
@@
Our ablations revealed several key insights: (1) Unlike prior work, tiling orthographic views decreased performance (24.8% vs. 31.2%); (2) Removing the LLM significantly reduced performance (20.0% vs. 31.2%), likely due to its strong priors and its role in conditioning consistent generation across views; (3) Directly bypassing the instruction to the Image Generator decreased performance by 9.5%, suggesting the LLM’s image token outputs preserve crucial task information. These findings validate the importance of each component in OG-VLA.
5 Real Robot Experiments
@@
We perform real-world experiments on a Franka Emika Panda arm mounted on a tabletop, with a single front-facing camera, as shown in Figure 1. We collect 3–5 demos for 4 real-world tasks—22 demos in total—with human-annotated keyframes and a motion planner [40] to achieve annotated keyframes. We train both the baseline and our model on this dataset. For our model, we augment each keyframe with 10 SE(3) perturbations and finetune the Arnold-pretrained OG-VLA@30k checkpoint for another 10k iterations with a batch size of 64. For $\pi_{0}$-FAST [4] baseline, we use the provided pretrained checkpoint trained on 10k+ hours of robot data across various robot setups with actions in both joint and end-effector control spaces. We finetune it on our dataset with joint angle actions for 30k iterations with a batch size of 32, as used in most of their pre-training and finetuning experiments [4, 22]. During inference, the model predicts an action chunk of 10 actions; we execute the last action in the sequence for faster execution.
@@
Table 2: Real world success rate (%). Success rates are reported for 4 real world tasks and novel pose, novel object, and novel scene test splits, averaged over 10 episodes each.
Quantitative Results
We report our results in Table 2. Each test set success rate is averaged over 10 episodes. For novel object variation, we use unseen colored and shaped objects. For novel scene variation, we introduce distractors and change lighting, background, and table appearance. Please refer to Appendix D for training and test scene details. Our results show that OG-VLA can adapt to new tasks with only 3–5 demonstrations and generalize well to novel poses, objects, and scenes. Although we attempted to compare with $\pi_{0}$-FAST, it failed at all tasks, learning only to reach the block with 30% success. We hypothesize that this is due to the small number of demonstrations and lack of support for SE(3) data augmentation without our 3D representation.
@@
6 Conclusion
We introduced OG-VLA, a novel architecture and learning framework that combines the generalization strengths of Vision-Language-Action (VLA) models with the robustness of 3D-aware keyframe policies for robotic manipulation. By leveraging foundation models in language and vision, OG-VLA improves generalization to unseen instructions, objects, and scenes while maintaining precise control. Our approach ensures input-view invariance by projecting multi-view RGBD observations into a canonical point cloud representation, which is processed through a vision backbone, an LLM, and an image diffusion model to generate actions as images. OG-VLA achieves state-of-the-art performance on robotic manipulation tasks, particularly on scene and object generalization tests. Moreover, OG-VLA can adapt to real-world tasks in 3-5 demonstrations with strong generalization to unseen objects and scenes. These results highlight the effectiveness of integrating pretrained visual and text priors into a structured 3D-aware framework for robot learning. Future work will explore extending OG-VLA to complex long-horizon tasks, external data augmentation and co-training techniques.
7 Limitations
OG-VLA has several strengths, but is not free of limitation. One challenge arises from the reliance on orthographic canonical views, which, while effective in many scenarios, can struggle in situations with severe occlusions, such as when multiple objects are stacked on a shelf against a wall. Occlusions may lead to partial or incorrect scene representations, potentially affecting the downstream task performance. Moreover, keyframe-based policies cover a broad range of tasks, however, are not able to solve several dynamic tasks, such as object tossing, or pressing with a desired force. Another consideration is the computational cost associated with OG-VLA. The current training procedure is computationally intensive, requiring substantial time and resources. We aim to address this by optimizing the model architecture, and exploring strategies such as distillation or parameter-efficient fine-tuning to accelerate learning while maintaining performance. Markovian policy learning also makes long horizon tasks quite challenging where several keyframes need to be predicted.
Real World Limitations
Single camera input makes some of the orthographic views redundant, uninformative or even noisy, which affects downstream performance. In this case, even if the predictions are correct in a few views, the noisy predictions affect the eventually decoded keyframe.
@@ A Tasks
Open Drawer 25, 50, 75, 100 (%) $\pm$10554
Close Drawer 0, 25, 50, 75 (%) $\pm$10671
Open Cabinet 25, 50, 75, 100 (%) $\pm$10319
Close Cabinet 0, 25, 50, 75 (%) $\pm$10478
Pour Water 25, 50, 75, 100 (%) $\pm$10312
Transfer Water 20, 40, 60, 80 (%) $\pm$10259
Table 3: Overview of the 8 tasks in ARNOLD. Each task features 4 goal states specified by human language, one of which is reserved for novel state evaluation and the other three are seen in the training dataset. The task is considered successful when the object state remains in the success range for two seconds. For Transfer Water, an additional condition of only less than 10% spillage of the original amount of water in the cup is imposed.
@@
Table 4: Examples of instruction templates used for the tasks.
@@ B.1 OG-VLA Prompt and Response
B.2 Text Action model Prompt and Response
@@ C Detailed Ablation Results and Analysis
C.1 Action Prediction Ablations
@@
1) Text Action: is an architecture variant where the LLM also produces gripper’s next pose in the form of text akin to prior VLA models [41]. In general, the raw output sequence from the LLM can be of slightly different format and contain additional text (e.g. text like the next robot action is: ), so long as it contains the information above. We apply regex parsing to extract from the text the gripper position $p^{text}$, orientation $\omega^{text}$. The prompt used for this ablation is shown in Appendix B.2.
Predicting actions as text without any visual reasoning on the output end of the model results in an imprecise action prediction model. Li et al. [41] have shown this architecture to work when training with large-scale robot datasets. However, we find that with a small robotics dataset, it is hard for VLAs to learn precise control via direct text token prediction.
2) Action Tokens: is an architecture variant where the LLM produces special action tokens added to the LLM vocabulary, such as $[trans0]$ or $[rot0]$ for translation and rotation modalities, akin to that for the image modality. This architecture is closer to works that perform action tokenization [3]. We decode the hidden state vectors for these tokens using additional MLP decoders to predict 3D translation and rotation vectors.
C.2 Image Generation Mode
@@
Table 5: Success rates for image generation modes for each task. We show that generating actions with reconstruction or with faded reconstructions work better than that without reconstruction.
We report success rates across tasks and test splits in Table 5. Generation without reconstruction of the scene performs the worst of all generation modes. This may be due to forcing the ImageGenerator to unlearn it’s prior of generating natural images and forcing it to only predict Gaussian distributions. We also observe instability in training this version, as sometimes the ImageGenerator would collapse to produce completely black or noisy images and stop generating actions on them. Generation with reconstruction and that with faded reconstruction did not show consistent difference over all test splits. Generation with reconstruction learns a policy that’s better by 4.4% on Novel Pose split and 2.3% Novel State split. Generation with faded reconstruction performs better on Novel Object and Novel State splits by 3.3% and 3%. Therefore, we conclude that these model variants have similar performance, and generation with faded reconstruction did not work better as we had hypothesized due to its balance between forcing the model to focus less on scene reconstruction and more on action prediction. Therefore, we report generation with reconstruction as our final method.
C.3 Model Ablations
@@
Table 6: Model ablation results for each task. We ablate components of OG-VLA to justify our design choices such as using tiled views, and the contribution of the LLM in the pipeline.
We present ablations on OG-VLA’s design choices in Table 6. The first result shows the effect of tiling the 4 orthographic views instead of feeding them to the ImageGenerator in batch of 4 as also explored in Genima [25]. For tiling, we stack the 4 views as in 2D array, following prior work. Tiling did not improve the performance for our model as opposed to results reported in the prior works. This may have occurred because the prior work did not have an LLM in the pipeline, so tiling becomes necessary for modeling interactions between views to generate multi-view consistent predictions. However, due to the LLM in OG-VLAconditioning generation in all views, there’s sufficient interaction and reasoning between views through the predicted image tokens $\left(t^{i}_{a},\quad i\in\{1,\dots,4\}\right)$.
Finally, we study the effect of bypassing the prompt and instruction into the ImageGenerator, shown in the last section of Table 6. The performance drop suggests that some instruction information may have been lost in image tokens output from the LLM, therefore it is beneficial to provide that information separately to the ImageGenerator for more accurate action prediction.

### v2

@@ IV Simulation Experiments
[…]
@@
Table I: Success rate on ARNOLD [gong2023arnold]. Success rates for 8 tasks and 4 test splits are shown for 2 baseline models and our model at specified training iterations (30k and 100k). The first row for each model is the Novel Pose split (the Test set in Arnold). The numbers in bold represent the best system for each task and test split.
IV-A Benchmarks and Datasets
We evaluate our method on the Arnold [gong2023arnold] and Colosseum [2024colosseum] benchmarks, which test language-grounded robot task learning in realistic 3D scenes with emphasis on testing generalization. Arnold uses five input cameras (front, base, left, wrist top, wrist bottom) and includes eight language-conditioned tasks (see Table I), each with four generalization test splits: (1) Novel Pose (held-out object/robot placements), (2) Novel Object (unseen objects), (3) Novel Scene (unseen scenes with seen objects), and (4) Novel State (unseen goal states). Arnold tasks follow a two-keyframe format—grasping and manipulating (e.g., pull drawer to 50% open)—and do not require gripper state prediction. These tasks demand both object pose and free space reasoning. Colosseum has four cameras (front, left shoulder, right shoulder, wrist) and features 20 language-conditioned tabletop tasks (e.g., close box, empty dishwasher) with 2–13 gripper keyframes (average 6), requiring gripper open/close prediction. It evaluates generalization via the all perturbation test set, which simultaneously alters object/table/background appearance, lighting, camera pose, and adds distractors.
@@ IV-A Benchmarks and Datasets
Each Arnold test split includes 20 episodes; Colosseum uses 25. We evaluate each Arnold model over 3 runs to account for simulator’s motion planner noise during keypoint execution, and we report means and standard deviations. Baseline models were trained by the respective benchmark authors. We re-run PerAct [shridhar2022peract] evaluation for 3 runs on Arnold for fair comparison, and report 6D-CLIPort [zheng2022vlmbench] results from Arnold [gong2023arnold] due to lack of checkpoint release.
For state-of-the-art (SotA) VLA comparisons, we include $\pi_{0}$-FAST [pertsch2025pi0fast] and $\pi_{0.5}$ [intelligence2025pi05visionlanguageactionmodelopenworld] baselines, reported in Table II. We compare single task VLA policies due to compute constraints. We finetune provided checkpoints pretrained on 10k+ hours of data across various robot setups with actions in both joint and end-effector continuous control spaces. All models are trained and evaluated on Pickup Object task with 600 training and 20 test episodes, batch size 64, and 30k iterations. Each $\pi$ model was finetuned on 4xL40 GPUs (LoRA: 0.5 days; Full: 2 days). At inference, we execute first half of the predicted action chunk (5 actions) for temporal consistency, inferring 80 times per episode to allow 400 environment steps.
IV-B Simulation Results
IV-B1 Arnold
@@
We present our model results in Table I. OG-VLA outperforms the baseline (PerAct) for most of the tasks’ Novel Pose split (seen objects and scenes), with a task-averaged relative improvement of 10.8%. We present our result at 30k and 100k iterations. OG-VLA improves significantly across generalization splits (Novel Object, Scene, and State), with relative increases of 20.0% and 46.5% respectively at 30k and 100k iterations in overall success rates. This demonstrates the effectiveness of our method in adapting pretrained visual and textual priors to robotics applications.
Model $\rightarrow$ $\pi_{0}$-FAST $\pi_{0}$-FAST $\pi_{0.5}$ OG-VLA
(LoRA) (Full) (Full) (LoRA)
Pickup Object 25.0 35.0 0.0 95.0
   - Novel Object 25.0 50.0 5.0 90.0
   - Novel Scene 20.0 35.0 0.0 80.0
   - Novel State 10.0 30.0 0.0    0.0
Table II: Comparison with VLAs (%). Success rates are reported for Pickup Object Arnold task on Novel Pose, Novel Object, Novel Scene and Novel State test splits, averaged over 20 episodes each.
Comparison with VLAs
SotA VLA baselines significantly underperform OG-VLA even on the simplest Arnold task (Table II). $\pi_{0}$-FAST learns the task, likely due to Arnold data’s action frequency being more suitable for its discrete frequency-space action tokenizer, unlike for the flow-matching decoder used for $\pi_{0.5}$. $\pi_{0}$-FAST shows promise on Novel State, reflecting its preserved language understanding, whereas OG-VLA struggles comparatively on this task, likely due to the lack of similar pretraining. Co-training with external datasets can improve OG-VLA, as show in prior works [yuan2024robopoint]. These results confirm that injecting 3D understanding in VLAs significantly improves them while preserving their generalization capabilities.
[…]
$\pi_{0}$-FAST (LoRA) 0.4 80 102.5 (42.1)
$\pi_{0}$-FAST 0.4 80 103.1 (39.1)
$\pi_{0.5}$ 0.2 80 70.3
PerAct 0.2    2    1.5
OG-VLA (LoRA) 4.5    2 10.2
Table III: Model Latency (seconds). Reported for per inference step, number of inference steps required per episode by each model, and finally overall episode latency with simulation execution for Pickup Object Arnold task, averaged over 20 episodes. indicates model inference time reported only on successful episodes.
Model Latency
Table III reports inference latency for both the baseline models and OG-VLA. While per-step inference latency is higher for OG-VLA due to multiple encoder–decoder modules and data processing involved during prediction, our model operates as a keyframe policy and thus requires only a few inference steps per episode, similar to other 3D keyframe policies [shridhar2022peract, goyal2023rvt]. As a result, overall episode inference time is much lower than VLA baselines, since the motion planner adds negligible cost compared to continuous control predictions in VLAs. However, OG-VLA remains slower than PerAct. We defer this optimization to future work.
@@
Figure 3: Evaluation on the COLOSSEUM benchmark [2024colosseum]. Task-averaged success rate shows that OG-VLA outperforms all baselines on the hardest generalization test set (all perturbation).
IV-B2 Colosseum
@@
Figure 3 shows task-averaged success rate on all perturbation test set on the Colosseum benchmark. We compare with baselines reported in Colosseum, including R3M [r3m], MVP [mvp], 3DDA [3d_diffuser_actor], RVT [goyal2023rvt], and PerAct [shridhar2022peract]. OG-VLA improves upon the baselines by a relative increase of 45.8%. The absolute performance remains low however (10.5%). We observe that Colosseum tasks are much harder to learn due to longer sequence of keyframe predictions, leading to error accumulation for an imitation learning based method.
IV-C Design Choice Ablations
@@
We conducted extensive ablations on the Arnold benchmark to validate our design choices. We highlight the main findings in this section, with detailed results and analysis in Appendix on project website.
@@
Our ablations revealed several key insights: (1) Unlike prior work, tiling orthographic views decreased performance (24.8% vs. 31.2%); (2) Removing the LLM significantly reduced performance (20.0% vs. 31.2%), likely due to its strong priors and its role in conditioning consistent generation across views; (3) Directly bypassing the instruction to Image Generator decreased performance by 9.5%, suggesting that LLM’s image token outputs preserve crucial task information. These findings validate the importance of each component in OG-VLA.
V Real Robot Experiments
@@
We perform real-world experiments on a Franka Emika Panda arm mounted on a tabletop, with a single front-facing camera, as shown in Figure 1. We collect 3–5 demos for 4 real-world tasks—22 demos in total—with human-annotated keyframes and a motion planner [zhu2022viola-deoxys] to achieve annotated keyframes. We train both the baseline and our model on this dataset. For our model, we augment each keyframe with 10 SE(3) perturbations and finetune the Arnold-pretrained OG-VLA@30k checkpoint for another 10k iterations with a batch size of 64. We use $\pi_{0}$-FAST Full finetuning as our baseline due to its strongest performance in simulation. We finetune it on our dataset with joint angle actions for 30k iterations with a batch size of 32, as used in most of their pre-training and finetuning experiments [pertsch2025pi0fast, black2024pi0]. During inference, the model predicts an action chunk of 10 actions; we execute the last action in the sequence for faster execution.
Quantitative Results
We report our results in Table IV. Each test set success rate is averaged over 10 episodes. For novel object variation, we use unseen colored and shaped objects. For novel scene variation, we introduce distractors and change lighting, background, and table appearance. Please refer to Appendix on project website for training and test scene details. Our results show that OG-VLA can adapt to new tasks with only 3–5 demonstrations and generalize well to novel poses, objects, and scenes. Although we attempted to compare with $\pi_{0}$-FAST, it failed at all tasks, learning only to reach the block with 30% success. We hypothesize that this is due to the small number of demonstrations and lack of support for SE(3) data augmentation without our 3D representation.
@@
Table IV: Real world success rate (%). Success rates are reported for 4 real world tasks and novel pose, novel object, and novel scene test splits, averaged over 10 episodes each.
@@
VI Conclusion and Limitations
We introduced OG-VLA, a novel architecture and learning framework that combines the generalization strengths of VLA models with the robustness of 3D-aware keyframe policies for robotic manipulation. By leveraging foundation models in language and vision, OG-VLA improves generalization to unseen instructions, objects, and scenes while maintaining precise control. Our approach ensures input-view invariance by projecting one or more RGBD observations into a canonical point cloud representation, which is processed through a vision backbone, an LLM, and an image diffusion model to generate actions as images. OG-VLA achieves state-of-the-art performance on robotic manipulation tasks, particularly on scene and object generalization tests. Moreover, OG-VLA can adapt to real-world tasks in 3-5 demonstrations with strong generalization to unseen objects and scenes. In practice, despite per-step inference being relatively slow, OG-VLA operates efficiently as a keyframe policy that predicts actions only at a low frequency, with intermediate motion handled by a motion planner. This allows fluid execution in real-world tasks even with a single camera, highlighting the flexibility of our framework. These results highlight the effectiveness of integrating pretrained visual and text priors into a structured 3D-aware framework for robot learning. Future work will explore extending OG-VLA to complex long-horizon tasks, external data augmentation, co-training techniques, and model optimization for faster inference.
OG-VLA has several strengths but also natural limitations. One challenge arises from the reliance on orthographic canonical views, which, while effective in many scenarios, can be affected by severe occlusions, such as multiple objects stacked on a shelf. Occlusions may lead to partial or incorrect scene representations, potentially affecting the downstream task performance. In practice, redundancy across multiple orthographic projections mitigates partial occlusions. Keyframe-based policies cover a broad range of tasks but are not suited for highly dynamic actions, such as object tossing or pressing with a desired force. Computational cost is another consideration, as training requires substantial time and resources; approaches such as distillation or parameter-efficient fine-tuning provide clear avenues for improvement. Markovian policy learning makes long-horizon tasks challenging, as multiple keyframes must be predicted. Finally, single-camera input may produce redundant or even noisy views, which can affect predictions; however, OG-VLA’s design ensures robustness to camera pose changes and allows future extensions with hierarchical or reactive controllers. Together, these limitations define clear directions for future research while demonstrating that OG-VLA provides a practical, generalizable, and robust approach to 3D-aware VLAs for robotic manipulation.
@@ A Tasks
Open Drawer 25, 50, 75, 100 (%) $\pm$10% 554
Close Drawer 0, 25, 50, 75 (%) $\pm$10% 671
Open Cabinet 25, 50, 75, 100 (%) $\pm$10% 319
Close Cabinet 0, 25, 50, 75 (%) $\pm$10% 478
Pour Water 25, 50, 75, 100 (%) $\pm$10% 312
Transfer Water 20, 40, 60, 80 (%) $\pm$10% 259
Table V: Overview of the 8 tasks in ARNOLD. Each task features 4 goal states specified by human language, one of which is reserved for novel state evaluation and the other three are seen in the training dataset. The task is considered successful when the object state remains in the success range for two seconds. For Transfer Water, an additional condition of only less than 10% spillage of the original amount of water in the cup is imposed.
@@
Table VI: Examples of instruction templates used for the tasks.
@@ B-A OG-VLA Prompt and Response
B-B Text Action model Prompt and Response
@@ C Detailed Ablation Results and Analysis
C-A Action Prediction Ablations
@@
1) Text Action: is an architecture variant where the LLM also produces gripper’s next pose in the form of text akin to prior VLA models [li2024llara]. In general, the raw output sequence from the LLM can be of slightly different format and contain additional text (e.g. text like the next robot action is: ), so long as it contains the information above. We apply regex parsing to extract from the text the gripper position $p^{text}$, orientation $\omega^{text}$. The prompt used for this ablation is shown in Appendix B-B.
Predicting actions as text without any visual reasoning on the output end of the model results in an imprecise action prediction model. [li2024llara] have shown this architecture to work when training with large-scale robot datasets. However, we find that with a small robotics dataset, it is hard for VLAs to learn precise control via direct text token prediction.
2) Action Tokens: is an architecture variant where the LLM produces special action tokens added to the LLM vocabulary, such as $[trans0]$ or $[rot0]$ for translation and rotation modalities, akin to that for the image modality. This architecture is closer to works that perform action tokenization [kim24openvla]. We decode the hidden state vectors for these tokens using additional MLP decoders to predict 3D translation and rotation vectors.
@@
Table VII: Success rates for image generation modes for each task. We show that generating actions with reconstruction or with faded reconstructions work better than that without reconstruction.
@@
Table VIII: Model ablation results for each task. We ablate components of OG-VLA to justify our design choices such as using tiled views, and the contribution of the LLM in the pipeline.
C-B Image Generation Mode